qwevery:

bar(s). Kaitlyn is interviewing applicants for an open job position in her department. During the interview process, applicants are required to type as much of a given handwritten document as possible in five minutes. The results for the three applicants are shown below. Applicant 1: 2 2/3 in 5 min Applicant 2: 3 1/4 in 5 min Applicant 3: 2 5/6 in 5 min Calculate each applicant's unit rate in pages typed per hour. Applicant 1's unit rate was _____ pages per hour. Applicant 2's unit rate was____ pages per hour. Applicant 3's unit rate was____ more pages per hour than applicant 1's unit rate.

snowflake0531:

60 minutes in an hour, 5 times 12 is 60 And to find the number of pages per hour multiply each fraction by 12

snowflake0531:

For applicant 1\[2 \frac{ 2 }{ 3 } * 12\] For applicant 2 \[3\frac{ 1 }{ 4 } * 12\] For applicant 3 \[2 \frac{ 5 }{ 6 } * 12\]

snowflake0531:

And so to find the last blank space subtract applicant 3 with applicant 1

qwevery:

thank you

snowflake0531:

yw~